# PreCi: Pre-training and Continual Improvement of Humanoid Locomotion via Model-Assumption-based Regularization

### v1

@@ I Introduction
Figure 1: Our Pre-training and Continual Improvement (PreCi) framework achieves a forward velocity of 1.5 m/s while successfully traversing a whiteboard covered with poppy seeds, as well as diverse outdoor terrains, including hills, uneven surfaces, and sand.
@@
To mitigate this problem, we introduce a novel learning framework, PreCi: Pre-training and Continual Improvement of a humanoid control policy via model-assumption-based regularization. Unlike the previous two-stage learning approach [23], our method introduces an adaptive regularization term based on the model assumption violation in the fine-tuning stage. More specifically, our method regularizes the policy to match the model-based controller’s actions if the robot’s state aligns with the assumptions of the underlying model. This approach is especially useful when the fine-tuning task extends beyond the capabilities of the model-based controller, as it reduces regularization in states where the robot violates model assumptions.
We demonstrate PreCi on a full-sized humanoid robot, Digit, in both simulation and real-world environments. In the simulation, our method can traverse both uneven and sloped terrains while achieving the highest tracking performance compared to the baseline methods. On hardware, our method can achieve diverse locomotion tasks including fast forward walking and robust walking across various indoor and outdoor terrains. Specifically, PreCi achieves a forward walking speed of 1.5 m/s while maintaining robustness on a whiteboard covered with poppy seeds, as well as on outdoor sloped, uneven, and deformable sandy terrains.
@@ II-A Model-based Control for Humanoid Locomotion
Humanoid locomotion often relies on optimization algorithms that use mathematical models to capture the robot’s essential dynamics. One of the most widely used models for bipedal locomotion is the linear inverted pendulum model (LIPM) [1], which provides analytical solutions for the foot placement to achieve a desired CoM trajectory [2, 6]. However, this simplified model fails to capture detailed joint-level behaviors, often resulting in conservative or infeasible full-body motions. To address these limitations, researchers have proposed more accurate models considering the inertia of the robot, such as single rigid body models  [3] and centroidal dynamics  [26, 27]. These inertial-informed models enable the online planning of the contact location, force, and centroidal states and have achieved agile locomotion skills such as running [28] and jumping [29]. Recently, with advancements in computational power, it has become feasible to plan with kino-dynamics [7] or even full-body dynamics model [8, 30]. Despite their impressive performance, model-based locomotion methods typically require accurate robot dynamics, well-informed environment setup, and a detailed specification of the gait sequence, which causes significant manual effort from experts.
@@ II-B Learning-based Approach for Humanoid Locomotion
In recent years, there have been significant advancements in legged locomotion due to the emergence of deep reinforcement learning algorithms and the power of massively parallel simulation environments. These learning frameworks initially demonstrated impressive performance in quadrupedal locomotion by optimizing policies based on carefully designed rewards [13, 14, 15], imitating reference motions [31], or incorporating model-based controllers [18, 19, 20]. Following this success, researchers began actively investigating learning-based frameworks for humanoid locomotion, presenting inherent instability and complexity. Some research groups proposed novel architectures [9, 10] to derive actions from contextual information embedded in observation histories. Others explored the use of demonstration data to streamline the training process and mitigate the reward engineering bottleneck by leveraging model-based methods [32, 33] or state-only motion capture data [10, 34].
@@ III Background: Model-based Controller
The inverse dynamics controller solves linearized dynamics $M\bm{\ddot{q}}^{\text{des}}-J^{T}\bm{\lambda}-S^{T}\bm{\tau}=-C\bm{\dot{q}}-G$, where $M$ is mass matrix, $J$ is contact jacobian, $S$ is selection matrix, $C$ is Coriolis and centrifugal term, and $G$ is the gravity vector. $\bm{\tau}$ is the joint torque, and $\lambda$ is the contact force. The inverse dynamics controller separates the linearized dynamics into actuated and unactuated parts to eliminate the contact force and solve for the desired joint torque. A passivity-based feedback controller is applied for stabilization [45] as feedback.
@@ IV Background: Imitating and Finetuning Model-based Controller
$L_{DAgger}=\sum_{\left(\bm{s},\bm{a^{E}}\right)\in\mathcal{D}}{\left\|\bm{a}^{E}-\mu(\bm{s})\right\|^{2}},$ (4)
where $\mathcal{D}$ is the aggregated data buffer, $\bm{s}$ and $\bm{a}^{E}=\pi^{E}(\bm{s})$ are the current robot state and the action of the given model-based controller, and $\mu$ is the mean network of the policy. This imitation gives us a pre-trained policy $\pi$ that behaves similarly to the model-based controller but also is learnable.
@@
V PreCi: Pre-training and Continual Improvement
In this section, we introduce a novel framework of Pre-training and Continual Improvement (PreCi), which is designed to improve the performance of the given model-based controller via pre-training and continual improvement. While our work is based on IFM [23], IFM often suffers from catastrophic forgetting, which results in poor performance in the real-world environment (Section V-A). To overcome this issue, we design a regularized loss that further distillates knowledge from the expert controller during fine-tuning (Section V-B). In addition, we introduce a model-assumption-based regularization (MAR) to determine the reliability of the expert controller (Section V-C). This allows us to reject unreliable data from MBC and thus improves the performance of the fine-tuned policy. Finally, we will describe a few implementation details, such as reward functions or Lipschitz Continuity Penalty (Section V-D).
@@ V-A Motivation: Catastrophic Forgetting of Motion Style
During RL-based fine-tuning, IFM often stabilizes its torso movement by adjusting swing leg abduction inward and reducing lateral momentum as shown in Fig. 2. However, this overfitted adaptation in a training environment can lead to potential instability due to the reduced size of the contact area, and this problem will be further exacerbated when the policy is deployed on hardware with a significant sim-to-real gap. While domain randomization may help mitigate this issue, its performance improvement is marginal and requires extensive manual tuning through trial and error.
@@
Figure 2: Motion forgetting example. Unlike MBC, IFM is trained to narrow its foot in the swing phase to optimize the tracking reward in the lateral direction.
@@ V-B Improved Fine-tuning with Regularization
$L_{FullReg}(\theta,\sigma)=\mathbb{E}\left[\ell_{PPO}(\theta,\sigma)+w\|\mathbf{a}^{E}-\mu_{\theta}(\mathbf{s})\|_{2}^{2}]\right],$ (5)
where $\mathbf{a}^{E}$ is the expert action of the given state $\mathbf{s}$, $\ell_{PPO}$ is the original PPO loss [47], and $w$ is a weight for the regularization loss. The policy is modeled as Gaussian distribution with the mean $\mu_{\theta}(\mathbf{s})$ and learnable standard deviation $\sigma$.
However, this straightforward regularization can result in a suboptimal policy because the regularization term limits the policy improvement that could be achieved by RL. In the extreme case of $\lambda\to\infty$, the learning framework simply replicates the expert’s behaviors without any improvements. Ideally, the weights should be dynamically adjusted, increasing the importance of the regularization term when the expert shows stable performance and vice versa.
@@ V-C Model-Assumption-based Regularization
$L_{PreCi}(\theta,\sigma)=\mathbb{E}\left[\ell_{PPO}(\theta,\sigma)+w(\mathbf{s})\|\mathbf{a}^{E}-\mu_{\theta}(\mathbf{s})\|_{2}^{2}\right],$ (6)
where $w:\mathbf{s}\rightarrow\left[0,\infty\right)$ is the weighting function given as follows with the tunable smoothing parameter $\sigma$ and its coefficient $w$:
$w(\mathbf{s})=we^{\frac{-\dot{z}^{2}}{\sigma}},$ (7)
This particular choice of the weighting function is based on the observation that the model assumption can be violated even when the controller performs successfully. By smoothly transitioning the weight to zero as violations increase, it dynamically adjusts the importance of expert knowledge for each sample by evaluating it against the given assumption. As a result, when a state deviates from the desirable state due to various factors, such as terrain changes or high command velocities, our framework automatically adjusts the weight of the regularization term for the corresponding states and achieves better performance. Moreover, our approach provides a more direct and intuitive adjustment of the loss function compared to the indirect tuning of reward functions [9, 22].
@@ V-D Implementation Details
Lin Vel Track $\exp(-\|\mathbf{v}_{xy}-\mathbf{v}_{xy}^{cmd}\|_{2}^{2}/\sigma_{xy})$ $1.2$
Ang Vel Track $\exp(-(\omega_{z}-\omega_{z}^{cmd})^{2}/\sigma_{\omega})$ $1.1$
Torque Penalty $\|\mathbf{\tau}\|_{2}^{2}$ $-4\times 10^{-6}$
Base Motion $\|\omega_{xy}\|_{2}^{2}$ $-0.6$
@@
$\min_{\theta,\sigma}\quad L_{PreCi}(\theta,\sigma)$ (8)
[…]
where L is the Lipschitz constant. Following the simplification process in [49], the given optimization is formulated as follows:
$\min_{\theta,\sigma}\quad L_{PreCi}(\theta,\sigma)+\alpha\mathbb{E}\left[\left\|\nabla_{\mathbf{s}}\mu_{\theta}(\mathbf{s})\right\|^{2}\right]$ (9)
where $\alpha$ is a tunable variable.
@@ VI Experimental Results
We design simulation tests and hardware experiments to investigate the following questions: (1) Can PreCi learn an effective policy in the training environments? (2) Can PreCi show robust performance in sim-to-sim and sim-to-real transfer scenarios compared to the baseline methods? (3) Can MAR dynamically adjust the sample weights based on the model-based assumption violation?
@@ VI-A1 Baselines
We consider the following baselines for our simulation tests and hardware experiments.
@@
  • IFM: IFM [23] learns a policy by imitating the given MPC and fine-tuning it. We employ the same reward configuration and Lipschitz penalty for smoothness.
@@
  • PureRL: This baseline is trained only with reinforcement learning without any pre-training or regularization. Its reward configuration is similar to [9] with 15 terms. However, we leverage a swing foot trajectory from an ALIP foot placement controller and add a reward for foot air time.
@@ VI-A2 Humanoid
As a robotic platform, we employ the Digit provided by Agility Robotics. This full-size humanoid robot has 20 motors, 8 motors in its upper body, and 12 motors in its lower body control. Also, Digit has 10 passive joints including spring joints in its shin and heel.
@@ VI-B Training Environment Experiments
Figure 4: Sloped (top) and uneven (bottom) terrains used for the robustness tests in MuJoCo.
Figure 5: Terrain level reached by each controller in the MuJoCo robustness test. PreCi successfully traverses all terrains within the time limit. IFM stumbles over its own feet on uneven terrain. FullReg fails to complete the final terrain due to poor tracking performance.
We evaluate the robustness of the controllers in MuJoCo on sloped and uneven terrains (Fig. 4). The robot is commanded to move forward at 0.6 m/s while adjusting its orientation to maintain a forward direction. As the robot moves forward, it traverses a sequence of terrains with increasing difficulty. The test is conducted with five different random seeds, and each controller runs until the robot either falls or reaches the time limit of 130 seconds.
We measure the terrain level reached by each controller shown in Fig. 5. Overall, PreCi (ours) and PureRL show the best performance among all baselines. While IFM reaches near the final level on sloped terrain, it often fails on uneven terrain due to unexpected foot contacts. These contacts cause severe lateral and angular disturbances, eventually leading the robot to trip over its own foot and fall. In contrast, PreCi robustly handles such disturbances as a result of its pre-trained motion and consistently reaches the final terrain level in both terrains. FullReg also demonstrates this level of robustness compared to IFM, but its tracking performance degrades under model-assumption violations, causing it to time out before reaching the goal. A more detailed discussion of this is provided in Section VI-E. PureRL also reaches the goal within the time limit on both terrains. However, the policy is overfitted to the simulation environment, resulting in degraded motion quality, which may lead to poor performance in both sim-to-sim and sim-to-real transfer. We discuss this further in Sections VI-C and VI-D.

### v2

# PPF: Pre-training and Preservative Fine-tuning of Humanoid Locomotion via Model-Assumption-based Regularization

@@ I Introduction
Figure 1: Our Pre-training and Preservative Fine-tuning (PPF) framework achieves a forward velocity of 1.5 m/s while successfully traversing a whiteboard covered with poppy seeds or olive oil, as well as diverse outdoor terrains, including hills, uneven surfaces, and sand.
@@
To mitigate this problem, we introduce a novel learning framework, PPF: Pre-training and Preservative Fine-tuning of a humanoid control policy via model-assumption-based regularization. Unlike the previous two-stage learning approach [23], our method introduces an adaptive regularization term based on the model assumption violation in the fine-tuning stage. More specifically, our method regularizes the policy to match the model-based controller’s actions if the robot’s state aligns with the assumptions of the underlying model. This approach is especially useful when the fine-tuning task extends beyond the capabilities of the model-based controller. For example, the ALIP model [26] assumes a constant body height and zero vertical velocity. When these assumptions are violated, such as during high-speed motion or traversal over rough terrain, the regularization weight is reduced, allowing the policy to improve beyond the limitations of the model.
We demonstrate PPF on a full-sized humanoid robot, Digit, in both simulation and real-world environments. In the simulation, our method can traverse both uneven and sloped terrains while achieving the highest tracking performance compared to the baseline methods. On hardware, our method can achieve diverse locomotion tasks including fast forward walking and robust walking across various indoor and outdoor terrains. Specifically, PPF achieves a forward walking speed of 1.5 m/s while maintaining robustness on a whiteboard covered with poppy seeds or olive oil, as well as on outdoor sloped, uneven, and deformable sandy terrains.
@@ II-A Model-based Control for Humanoid Locomotion
Humanoid locomotion often relies on optimization algorithms that use mathematical models to capture the robot’s essential dynamics. One of the most widely used models for bipedal locomotion is the linear inverted pendulum model (LIPM) [1], which provides analytical solutions for the foot placement to achieve a desired CoM trajectory [2, 6]. However, this simplified model fails to capture detailed joint-level behaviors, often resulting in conservative or infeasible full-body motions. To address these limitations, researchers have proposed more accurate models considering the inertia of the robot, such as single rigid body models  [3] and centroidal dynamics  [27, 28]. These inertial-informed models enable the online planning of the contact location, force, and centroidal states and have achieved agile locomotion skills such as running [29] and jumping [30]. Recently, with advancements in computational power, it has become feasible to plan with kino-dynamics [7] or even full-body dynamics model [8, 31]. Despite their impressive performance, model-based locomotion methods typically require accurate robot dynamics, well-informed environment setup, and a detailed specification of the gait sequence to obtain this performance, which causes significant manual effort.
@@ II-B Learning-based Approach for Humanoid Locomotion
In recent years, there have been significant advancements in legged locomotion due to the emergence of deep reinforcement learning algorithms and the power of massively parallel simulation environments. These learning frameworks initially demonstrated impressive performance in quadrupedal locomotion by optimizing policies based on carefully designed rewards [13, 14, 15], imitating reference motions [17], or incorporating model-based controllers [18, 19, 20]. In particular, incorporating model-based methods enables learning-based methods to acquire periodic and symmetric gaits in a more interpretable way, mitigating the tedious reward shaping. The similar success on quadruped has been replicated on humanoid locomotion via learning-based frameworks. Some studies proposed novel architectures [9, 10] to derive actions from contextual information embedded in observation histories. Others explored the use of demonstration data to streamline the training and mitigate the reward engineering bottleneck by leveraging model-based methods [32, 33] or state-only motion capture data [10, 34].
@@ III Background: Model-based Controller
The inverse dynamics controller solves linearized dynamics $M\bm{\ddot{q}}^{\text{des}}-J^{T}\bm{\lambda}-S^{T}\bm{\tau}=-C\bm{\dot{q}}-G$, where $M$ is mass matrix, $J$ is contact jacobian, $S$ is selection matrix, $C$ is Coriolis and centrifugal term, and $G$ is the gravity vector. $\bm{\tau}$ is the joint torque, and $\bm{\lambda}$ is the contact force. The inverse dynamics controller separates the linearized dynamics into actuated and unactuated parts to eliminate the contact force and solve for the desired joint torque. A passivity-based feedback controller is applied for stabilization [44] as feedback.
@@ IV Background: Imitating and Finetuning Model-based Controller
$L_{DAgger}=\sum_{\left(\bm{s},\bm{a^{E}}\right)\in\mathcal{D}}{\left\|\bm{a}^{E}-\mu_{\theta}(\bm{s})\right\|^{2}},$ (4)
where $\mathcal{D}$ is the aggregated data buffer, $\bm{s}$ and $\bm{a}^{E}$ denote the sampled robot state and the corresponding action from the expert model-based controller, where $E$ stands for the “expert”. $\mu_{\theta}$ denotes the neural network policy trained via supervised learning. While deterministic here, it later serves as the mean of the Gaussian policy during RL fine-tuning. This imitation gives us a pre-trained policy $\pi$ that behaves similarly to the model-based controller but also is learnable.
@@
V PPF: Pre-training and Preservative Fine-tuning
In this section, we introduce Pre-training and Preservative Fine-tuning (PPF), a novel framework designed to enhance the performance of a given model-based controller by first distilling its knowledge into a learnable neural network, and then fine-tuning it using RL with model-assumption-based regularization. While our work is based on Imitating and Finetuning Model Predictive Control (IFM) [23], IFM often suffers from catastrophic forgetting, which results in poor performance in the real-world environment (Section V-A). To overcome this issue, we design a regularized loss that further distills knowledge from the expert controller during fine-tuning (Section V-B). In addition, we introduce a model-assumption-based regularization (MAR) to determine the reliability of the expert controller (Section V-C). This allows us to reject unreliable data from MBC and thus improves the performance of the fine-tuned policy. Finally, we will describe a few implementation details, such as reward functions or Lipschitz Continuity Penalty (Section V-D).
@@ V-A Motivation: Catastrophic Forgetting of Motion Style
During RL-based fine-tuning, IFM often stabilizes the pelvis lateral movement by initially swinging the leg inward and stepping with narrower foot placement as shown in Fig. 2. However, this overfitted adaptation in a training environment can lead to potential instability, and this problem will be further exacerbated when the policy is deployed on hardware with a significant sim-to-real gap. For instance, frequent foot collisions were observed when deploying the IFM in both the AR-sim and on hardware. While domain randomization may help mitigate this issue, its performance improvement is marginal and requires extensive manual tuning through trial and error.
@@
Figure 2: Motion forgetting example. Unlike MBC, IFM is trained to swing its foot inwards initially and step with narrower foot placement, optimizing for lateral tracking accuracy and reduced energy consumption.
@@ V-B Improved Fine-tuning with Regularization
$L_{FullReg}(\theta,\sigma)=L_{PPO}(\theta,\sigma)+w\mathbb{E_{(\mathbf{s},\mathbf{a^{E}})\sim\mathcal{D}}}\left[\|\mathbf{a}^{E}-\mu_{\theta}(\mathbf{s})\|_{2}^{2}\right],$ (5)
where $\mathbf{a}^{E}$ is the expert action of the given state $\mathbf{s}$, $L_{PPO}$ is the PPO loss, and $w$ is a weight for the regularization loss. The policy is modeled as Gaussian distribution with the mean $\mu_{\theta}(\mathbf{s})$ and learnable standard deviation $\sigma$.
However, this straightforward regularization can result in a suboptimal policy because the regularization term limits the policy improvement that could be achieved by RL. In the extreme case of $w\to\infty$, the learning framework simply replicates the expert’s behaviors without any improvements. Ideally, the weights should be dynamically adjusted, increasing the importance of the regularization term when the expert shows stable performance and vice versa.
@@ V-C Model-Assumption-based Regularization
$L_{PPF}(\theta,\sigma)=L_{PPO}(\theta,\sigma)+\mathbb{E_{(\mathbf{s},\mathbf{a^{E}})\sim\mathcal{D}}}\left[w(\mathbf{s})\|\mathbf{a}^{E}-\mu_{\theta}(\mathbf{s})\|_{2}^{2}\right],$ (6)
where $w:\mathbf{s}\rightarrow\left[0,\infty\right)$ is the weighting function given as follows with the smoothing parameter $\delta$ and coefficient $w_{0}$:
$w(\mathbf{s})=w_{0}e^{\frac{-{\dot{z}(\mathbf{s})}^{2}}{\delta}},$ (7)
In our case, $w_{0}=5$ and $\delta=0.0159$ to preserve the motion in reliable states. This particular choice of the weighting function is based on the observation that the model assumption can be violated even when the controller performs successfully. By smoothly transitioning the weight to zero as violations increase, it dynamically adjusts the importance of expert knowledge for each sample by evaluating it against the given assumption. As a result, when a state deviates from the desirable state due to various factors, such as terrain changes or high command velocities, our framework automatically adjusts the weight of the regularization term for the corresponding states and achieves better performance. Moreover, our approach provides a more direct and intuitive adjustment of the loss function compared to the indirect tuning of reward functions [9, 22].
@@ V-D Implementation Details
Observation and action space. The observation space includes root linear and angular velocities, projected gravity, user command, lower-body joint positions and velocities, gait phase and domain, previous action, and the history of lower-body joint positions and velocities. The action space consists of target joint positions and velocities for the lower body.
@@
Lin Vel Track $\exp(-\|\mathbf{v}_{xy}-\mathbf{v}_{xy}^{cmd}\|_{2}^{2}/\delta_{xy})$ $1.2$
Ang Vel Track $\exp(-(\omega_{z}-\omega_{z}^{cmd})^{2}/\delta_{\omega})$ $1.1$
Torque Penalty $\|\mathbf{\tau}\|_{2}^{2}$ $4\times 10^{-6}$
Base Motion $\|\omega_{xy}\|_{2}^{2}$ 0.6-0.6
@@
Training Time and Data Size The model is trained using a GeForce RTX 4090 GPU and an Intel i9-14900K CPU. We use MuJoCo [48] for the simulator. The DAgger stage runs for 380 iterations, corresponding to 9.12 million samples, and takes approximately 5 hours and 26 minutes. The fine-tuning stage completes 10,000 iterations, generating approximately 400 million samples over 31 hours and 33 minutes.
@@
$\min_{\theta,\sigma}\quad L_{PPF}(\theta,\sigma)$ (8)
$\text{s.t.}\quad\max_{\mathbf{s}}\left\|\nabla_{\mathbf{s}}\mu_{\theta}(\mathbf{s})\right\|^{2}\leq K^{2},$
where $K$ is the Lipschitz constant. Following the simplification process in [49], it is formulated as follows:
$\min_{\theta,\sigma}\quad L_{PPF}(\theta,\sigma)+\alpha\mathbb{E}\left[\left\|\nabla_{\mathbf{s}}\mu_{\theta}(\mathbf{s})\right\|^{2}\right]$ (9)
where $\alpha$ is a tunable variable set to $\alpha=0.0001$ in our case.
@@ VI Experimental Results
We design simulation tests and hardware experiments to investigate the following questions: (1) Can PPF learn an effective policy in the training environments? (2) Can PPF show robust performance in sim-to-sim and sim-to-real transfer scenarios compared to the baseline methods? (3) Can MAR dynamically adjust the sample weights based on the model-based assumption violation?
@@ VI-A1 Baselines
We consider the following baselines for our simulation tests and hardware experiments. All neural network-based controllers use a multi-layer perceptron with hidden layers of sizes 512, 256, and 64.
@@
  • IFM: IFM [23] learns a policy by imitating the given MBC and fine-tuning it. We employ the same reward configuration and Lipschitz penalty.
@@
  • PureRL: PureRL is trained with reinforcement learning without any pre-training or regularization. Its reward configuration is similar to [9]. However, we leverage a swing foot trajectory from an ALIP foot placement controller, add rewards for foot air time and second-order action smoothing, and exclude the selected joint position penalty and target joint position smoothing.
@@ VI-A2 Humanoid
As a robotic platform, we employ the Digit provided by Agility Robotics. This has the same hardware configuration as described in [9].
@@ VI-B Training Environment Experiments
Figure 4: Testing terrains for the robustness tests in MuJoCo.
Figure 5: Terrain level reached by each controller in the MuJoCo robustness test. PPF successfully traverses all terrains within the time limit. IFM stumbles over its own feet on uneven terrain. FullReg fails to complete the final terrain due to poor tracking performance.
We evaluate the robustness of the controllers in MuJoCo on sloped and uneven terrains (Fig. 4). The robot is commanded to move forward at 0.6 m/s while adjusting its orientation to maintain a forward direction. As the robot moves forward, it traverses a sequence of terrains with increasing difficulty. The test is conducted with five different random seeds, and each controller runs until the robot either falls, reaches the time limit of 130 seconds, or reaches the end of the last level of the terrain (goal location).
We measure the terrain level reached by each controller shown in Fig. 5. Overall, PPF (ours) and PureRL show the best performance among all baselines. While IFM reaches near the final level on sloped terrain, it often fails on uneven terrain due to unexpected foot contacts. These contacts cause severe lateral and angular disturbances, eventually leading the robot to trip over its own foot and fall. In contrast, PPF robustly handles such disturbances as a result of its pre-trained motion and consistently reaches the final terrain level in both terrains. FullReg also demonstrates this level of robustness compared to IFM, but its tracking performance degrades under model-assumption violations, causing it to time out before reaching the goal. A more detailed discussion of this is provided in Section VI-E. PureRL also reaches the goal within the time limit on both terrains. However, the policy is overfitted to the simulation environment, resulting in degraded motion quality, which may lead to poor performance in both sim-to-sim and sim-to-real transfer. We discuss this further in Sections VI-C and VI-D.